# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

### v1

@@ 4 Method
The standard EKF strategy for online learning presented in Section 3.2 does not typically scale to large neural networks because of the memory cost is $O({D_{\bm{\theta}}}^{2})$ and the compute cost is $O({D_{\bm{\theta}}}^{3})$. A number of approaches are proposed to tackle this issue (see Appendix 6 for details). Our method leverages the frequentist perspective of filtering to derive low-rank updates and the Bayesian interpretation of filtering to make sequential decisions through the posterior predictive.
@@ 4.1 Algorithm
In this section, we sketch our algorithm. For further details and full derivation, see Appendix D.1.
@@
$\mathbb{E}[{\boldsymbol{\ell}}_{0}]$ $={\boldsymbol{\ell}}_{0|0},$ $\mathbb{E}[{\bm{h}}_{0}]$ $={\bm{h}}_{0|0},$ ${\rm Var}({\boldsymbol{\ell}}_{0})$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0},$ ${\rm Var}({\bm{h}}_{0})$ $=\mathbf{C}_{0}^{\intercal}\,\mathbf{C}_{0}.$ (2)
@@
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\$ (3)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables with dynamics covariance matrices $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$.
@@
Our method maintains an approximate belief state for the state-space model represented by the tuple $({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2})$, where ${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ is the BLUP estimate for model parameters in the last layer, ${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ is the BLUP estimate for model parameters in the hidden layers (feature extractor), $\hat{\boldsymbol{\Sigma}}_{t-1}^{1/2}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ is the Cholesky EVC factor for the covariance of the last layer, and $\mathbf{C}_{t}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ is the low-rank EVC factor for the hidden layers. Algorithm 1 shows a single step of HiLoFi.
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\mathbf{C}_{t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  // BLUP update step
15:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
16:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
17:  // EVC low-rank and Cholesky update step
18:  $\mathbf{C}_{t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
19:  $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\leftarrow{\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
@@
20:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\right)$
@@
Algorithm 1 Single step of HiLoFi for online learning. The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$; see Appendix C.2 for details.
The mean update step has the form ${\bm{\theta}}_{t|t}={\bm{\theta}}_{t|t-1}+\mathbf{K}_{t}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}={\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$ is the error or innovation term, and $\mathbf{K}_{t}$ is the Kalman gain matrix. This has roughly the same form as the EKF, except that $\mathbf{K}_{t}$ is computed separately for the last and hidden layers.
The EVC update however, is more complicated, since it requires the approximation of a block-diagonal covariance: the last layer block is full-rank, and the hidden layer blocks are low-rank. To maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers. The Cholesky factor is taken from a surrogate matrix that avoids computing a full ${\mathcal{M}_{{D_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ matrix by neglecting the artificial dynamics covariance $q_{{\boldsymbol{\ell}},t}$. The low-rank component for hidden layers is approximated via a two-step process: (i) a fast surrogate predicted covariance, and (ii) a rank ${d_{{\bm{h}}}}$ projection. Like the Cholesky factor, the surrogate covariance in (i) reduces the computational cost by ignoring the dynamics noise However, we retain some information from the dynamics noise to update the hidden layer parameters. Step (ii) enables a fast update rule.
After every update, the induced posterior predictive distribution is
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})={\cal N}\left({\bm{y}}_{t}\,|\,\tilde{\mathbf{L}}\,{\boldsymbol{\ell}}_{t|t-1}+\tilde{\mathbf{H}}\,{\bm{h}}_{t|t-1},\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\tilde{\mathbf{H}}^{\intercal}+\mathbf{R}_{t}\right),$ (4)
which is a proper density as long as $(\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\tilde{\mathbf{H}}^{\intercal}+\mathbf{R}_{t})\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. This can be easily ensured by taking $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. In (4), $\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\tilde{\mathbf{H}}^{\intercal}$ is the epistemic uncertainty and $\mathbf{R}_{t}$ is the aleatoric uncertainty.
Whenever the output layer or the input to the last layer is high-dimensional, we can perform a low-rank approximation to the last-layer to reduce computational costs. We call this variant LoLoFi. Its computational complexity is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. See Appendix E.2 for details.
A purely low-rank version of our method, which we call LRKF, is discussed in Appendix E.3. This method uses a single low rank covariance matrix approximation for all parameters in the neural network. In the experiments, we find that this approach has lower performance than HiLoFi, which is not surprising. It also has lower performance than LoLoFi (at least on the contextual bandit problem in section 5.2), presumably because modeling the last layer separately allows us to use a different subspace for the low rank approximation of the output layer weights, compared to use a shared subspace.
4.2 Theoretical properties
@@
where $\mathbf{E}_{{\bm{h}},{\rm surr}}=(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2})$, $\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}=\left\|\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\right\|_{\rm F}$, and $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are smallest $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
@@
Proposition 4.1, whose proof is in Appendix F.4, shows that the per-step error in the HiLoFi approximation is bounded by: (i) the neglected terms by each surrogate matrix, (ii) the low-rank approximation for the covariance of the hidden layers, and (iii) the cross-terms from the blocks in the off-diagonal. (Note, however, that it does not account for error due to linearization.) One could minimize the above lower bound by setting $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$. However, this may result in lower per-step performance. See Appendix G.1.2 for an example of the main sources of error.
@@ 5 Experiments
In this section, we experimentally evaluate our method. We start by considering a 1d example of "in-between" uncertainty, to qualitatively illustrate the behavior of the method. We then quantitatively compare its performance to other methods across the following sequential decision-making tasks:
@@
    MNIST for classification and bandits: This is an online high-dimensional classification problem, which we also study as a contextual bandit problem.
@@
For each of the above tasks, we assess the various methods on predictive performance and wall-clock time. All experiments were run on a TPU v4-8.
We consider the following variants of our framework. High-rank low-rank filter (HiLoFi): Our main method which uses full rank for last layer and low rank for hidden layers. Low-rank low-rank filter (LoLoFi): A version of our method that uses low-rank matrices for last and hidden layers. We only consider LoLoFi in task (II), due to the relatively large dimension of the last layer. Low-rank square-root Kalman Filter (LRKF): This is a special case of our method, described in Appendix E.3, that uses a single low-rank covariance for all the parameters. This serves to illustrate the gains from modeling the final layer separately.
As existing baselines, we consider the following. Diagonal plus low-rank precision (LoFi) [10]: A fully online method that models the precision matrix using a low-rank plus diagonal matrix, and uses the same linearization scheme as us. Gaussian processes (GP) [64]: A standard approach to modeling predictive uncertainty. We only include GPs in the low-dimensional tasks in (III) because of scalability limitations. Variational Bayesian last layer (VBLL) [7]: A partially Bayesian method that requires access to the full dataset (or a replay buffer), and performs multiple optimization steps per update. See Appendix E.4 for details. Online variational Bayesian last-layer (OnVBLL): A straightforward modification of the VBLL method with a first-in-first-out (FIFO) buffer and no regularization. Last-layer Laplace approximation with FIFO buffer (LLL) [13]: A method that uses a Laplace approximation on the last layer, with a FIFO replay buffer and multiple inner updates per time step. Unless otherwise stated, we train the parameters of the neural network using the AdamW optimization algorithm [44].
As an even simpler baseline, which does not model uncertainty, we consider Muon (muon) [32], which a special case of the Shampoo optimizer [26]. This is a quasi second-order optimization method that is scalable and shows strong empirical performance. We use the implementation in the Optax library [14]. Unlike the other methods, it only computes a point estimate, so we cannot use it for computing the posterior predictive. We use this method for the $\epsilon$-greedy bandit and as a choice of optimizer for LLL in Section 5.2.

### v2

@@ 4 Method
The standard EKF strategy for online learning presented in Section 3.2 does not scale to large neural networks because the memory cost is $O({D_{\bm{\theta}}}^{2})$ and the compute cost is $O({D_{\bm{\theta}}}^{3})$. A number of approaches have been proposed to tackle this issue (see Section 6 for details). Our method leverages the frequentist perspective of filtering to derive low-rank updates and the Bayesian interpretation of filtering to make sequential decisions through the posterior predictive.
@@ 4.1 Algorithm
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\\$ (8)
@@
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables with covariance matrices $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$, with $q_{{\boldsymbol{\ell}},t}\geq 0$ and $q_{{\bm{h}},t}\geq 0$. Our method maintains a belief state for the state-space model represented by ${\bm{b}}_{t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t})$, where ${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ is the estimate for model parameters in the last layer, ${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ is the estimate for model parameters in the hidden layers, $\hat{\boldsymbol{\Sigma}}_{t-1}^{1/2}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ is the Cholesky EVC factor for the covariance of the last layer, and $\mathbf{C}_{t}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ is the low-rank EVC factor for the hidden layers.
Updates.
Given ${\bm{b}}_{t-1}$, the updated ${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t})$ follows directly from the EKF equations which takes the form ${\bm{\theta}}_{t|t}={\bm{\theta}}_{t-1|t-1}+\mathbf{K}_{t}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}={\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$ is the error or innovation term, and $\mathbf{K}_{t}$ is the gain matrix. Because of the block-diagonal assumption, it can be shown that $\mathbf{K}_{t}$ is decoupled into last-layer-only and a hidden-layers-only submatrices.
Next, applying the EKF EVC update to ${\bm{b}}_{t-1}$ yields the dense matrix $\boldsymbol{\Sigma}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$ which couples the effects from the last and hidden layers. To obtain a block-diagonal matrix with components $(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t})$, we consider a two-stage approximation. The first stage replaces $\boldsymbol{\Sigma}_{t}$ with a surrogate covariance matrix $\tilde{\boldsymbol{\Sigma}}_{t}$ that avoids the interaction between the hidden layer parameters and the synthetic dynamics $q_{{\bm{h}},t}$; the second stage approximates the surrogate matrix with the best block-diagonal approximation whose first block (for last-layer parameters) are of rank $D_{{\boldsymbol{\ell}}}$} and the second block (for hidden-layer parameters) are rank-$d_{{\bm{h}}}$}. Finally, to maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers, i.e., $\hat{\boldsymbol{\Sigma}}_{t}^{1/2}={\rm diag}(\mathbf{C}_{t},\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2})$. In this sense, estimation of the factors for the last-layer and the hidden layer follows the sequence
$\boldsymbol{\Sigma}_{t}\to\tilde{\boldsymbol{\Sigma}}_{t}\to\hat{\boldsymbol{\Sigma}}_{t}\to\hat{\boldsymbol{\Sigma}}_{t}^{1/2}={\rm diag}(\mathbf{C}_{t},\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}).$
Algorithm 2 shows a single step of HiLoFi. The derivation is in Appendix D.1.
[…]
5:// predict step
[…]
9:// innovation (one-step-ahead error) and Cholesky innovation variance
[…]
12:// gain for hidden layers
[…]
15:// gain for last layer
[…]
18:// mean update step
[…]
21:// EVC low-rank and Cholesky update step
[…]
@@
24:Return ${\bm{b}}_{t}=\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t}\right)$ // updated belief
@@
Algorithm 2 Single update step of HiLoFi. The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$; see Appendix C.2 for details.
The following proposition shows the upper bound of the the per-step error incurred by HiLoFi.
@@
where $\mathbf{E}_{{\bm{h}},{\rm surr}}=(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2})$, $\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}=\|(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\,(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}\|_{\rm F}$, and $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are smallest $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
@@
Proposition 4.1 shows that the per-step error in the HiLoFi approximation is bounded by: (i) the neglected terms by each surrogate matrix, (ii) the low-rank approximation for the covariance of the hidden layers, and (iii) the cross-terms from the blocks in the off-diagonal Note, however, that the bound does not account for error due to linearization of the neural network. One could minimize the above lower bound by setting $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$. However, this may result in lower per-step performance. See Appendix G.1.2 for an example of the main sources of error. See Appendix E.4 for a proof.
Posterior predictive model.
After every update, the posterior predictive model used for sequential decision making is
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1})={\cal N}\left({\bm{y}}_{t}\mid{f}({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},{\bm{x}}_{t+1}),\;\;\tilde{\mathbf{L}}_{t+1}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t+1}^{\intercal}+\tilde{\mathbf{H}}_{t+1}\,\boldsymbol{\Sigma}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t+1}^{\intercal}+\mathbf{R}_{t+1}\right).$ (10)
Variants.
Whenever the output layer or the input to the last layer is high dimensional, we perform a low-rank approximation to the last-layer to reduce computational costs. We call this variant LoLoFi. Its computational complexity is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. See Appendix F.1 for details.
A purely low-rank version of our method, which we call LRKF, is discussed in Appendix F.2. This method uses a single low rank covariance matrix approximation for all parameters in the neural network. An error analysis of LRKF in the linear setting is presented in Appendix F.2.1. In the experiments, we find that this approach performs worse than that of HiLoFi, which is not surprising. It also has lower performance than LoLoFi (at least on the contextual bandit problem in section˜5.2), presumably because modeling the last layer separately allows us to use a different subspace for the low rank approximation of the output layer weights.
@@ 5 Experiments
In this section, we evaluate our method. First, we consider a one-dimensional example “in-between” uncertainty, to illustrate the behavior of the method. We then compare its performance to that of other methods across the following sequential decision-making tasks:
@@
    MNIST for classification as a bandit problem: This is an online high-dimensional classification problem, which we study as a contextual bandit problem.
@@
For each of the above tasks, we assess the various methods on predictive performance and wall-clock time. All experiments were run on a TPU v4-8. An additional experiment demonstrating the scalability of LRKF to multi-million-parameter neural networks for online classification is provided in Appendix G.2.
We consider the following variants of our framework. High-rank low-rank filter (HiLoFi): Our main method that uses full rank for last layer and low rank for hidden layers. Low-rank low-rank filter (LoLoFi): A version of our method that uses low-rank matrices for last and hidden layers. We only consider LoLoFi in task (II), due to the relatively large dimension of the last layer. Low-rank square-root Kalman Filter (LRKF): This is a special case of our method, described in Appendix F.2, that uses a single low-rank covariance for all the parameters. This illustrates the gains from modeling the final layer separately.
As existing baselines, we consider the following. Diagonal plus low-rank precision (LoFi) [8]: a fully online method that models the precision matrix using a low-rank plus diagonal matrix, and uses the same linearization scheme as us. Gaussian processes (GP) [62]: a standard approach to modeling predictive uncertainty. We only include GPs in the low-dimensional tasks in (III) because of scalability limitations. Variational Bayesian last layer (VBLL) [5]: a partially Bayesian method that requires access to the full dataset (or a replay buffer), and performs multiple optimization steps per update. See Appendix F.3 for details. Online variational Bayesian last-layer (OnVBLL): a straightforward modification of the VBLL method with a first-in-first-out (FIFO) buffer and no regularization. Last-layer Laplace approximation with FIFO buffer (LLL) [11]: a method that uses a Laplace approximation on the last layer, with a FIFO replay buffer and multiple inner updates per time step. Unless otherwise stated, we train the parameters of the neural network with the AdamW optimization algorithm [44].